|Project name||Main project page|
Previous entry Next entry
The IR mirror and QPD problem is resolved. Skipper reinstalled the mirror (inside the black box) and I realigned it,now it seems like, it works.
The QPD is more (enough) maneuverable inside the limit of ½ cm. The signals are not running crazily all over the place as before. After a short tweaking of x and y axis and the condenser they are easily overlapped at 0 amplitude. The signals look kinda noisy, skipper said, “He will check it up!” So I am moving to the next task now. More pics: http://www.evernote.com/pub/pranavrathi1/qpd